Without any stoppage, a person travels a certain distance at an average speed of 42 km/h, and with stoppages, he covers the same distance at an average speed of 28 km/hr. How many minutes per hour does he stop?

14 minutes
15 minutes
28 minutes
None of these

Kindly, tell me how to solve this question, please.

Thank you.
1
Expert's answer
2020-08-06T16:46:49-0400

What is the average speed? It is total distance over total time:


"v=\frac{D}{t}."

Thus, for 42 km/h and no stops we have


"42=\frac{D}{t_1}."

For 28 km/h it takes longer because certain time "t_s" is required to stop:


"28=\frac{D}{t_1+t_s}."

From the last equation:


"t_s=\frac{D}{28}-t_1."

Substitute "t_1"expressed from the equation with 42 km/h:


"t_s=\frac{D}{28}-\frac{D}{42}=\frac{D}{84}."

Compare this time with "t_1":


"t_1=\frac{D}{42},\\\space\\
\frac{t_s}{t_1}=\frac{1}{2},\\\space\\
t_s=0.5t_1."

Thus, stops last half of time that the traveler moves. They last 30 minutes per hour.